Find the absolute value of the complex number. 3i-2
Found 2 solutions by stanbon, joyofmath:
Answer by stanbon(48536)

absolute value of the complex number. 3i-2
|3i-2|= sqrt(3^2+2^2)=sqrt(13)
Cheers,
Stan H.

Answer by joyofmath(189)